OpenStudy (anonymous):

The amount P is invested at interest rate r compounded annually and grows to the amount A in two years. Find the interest rate. P=4000, A=5267.03

OpenStudy (anonymous):

ranga help

OpenStudy (anonymous):

Use the formula: \[A = P(1+\frac{ i }{ n })^{nt}\]

OpenStudy (anonymous):

annual compounding means n= 1 can you figure it out now?

OpenStudy (anonymous):

A=P(1+r)^2 is what my textbook has listed. What does i and n stand for?

OpenStudy (anonymous):

i is interest rate and n number of times interest is comunded

OpenStudy (anonymous):

Oh, okay. The equation I had down before it was 5267.03=4000(1+r)^2. Is that set up correctly (according to the textbook)? Because my teacher only wants us to solve it that way. Also, my main problem is solving the equation.

OpenStudy (anonymous):

yea its setup correctly

OpenStudy (anonymous):

you just need to isolate r by itself

OpenStudy (anonymous):

Okay. I tried dividing 5267.03 by 4000 initially, but then I got stuck with a very weird not-trinomial square thing after trying to isolate r. What's the right way to do it?

OpenStudy (anonymous):

i give you the first step divide both sides by 4k

OpenStudy (anonymous):

yea you got the first step right

OpenStudy (anonymous):

you should have something like this \[\frac{ 5267.03 }{ 4000 } = (1+r)^2\]

OpenStudy (anonymous):

Yeah, I have that. What would be the next step? I already checked to see if 5267.03 had a square root, but it didn't. So how do I continue?

OpenStudy (anonymous):

can you tell me the next step?

OpenStudy (anonymous):

you can just put it in your calculator

OpenStudy (anonymous):

if you isolate r you get \[r = \sqrt{\frac{ 5267.03 }{ 4000 }} - 1\] tell me what you get as your answer from your calculator

OpenStudy (anonymous):

Which part? The fraction?

OpenStudy (anonymous):

everything

OpenStudy (anonymous):

i set everything up for you so you just need to solve for r

OpenStudy (anonymous):

Thanks. Wait a second.

OpenStudy (anonymous):

I got 0.1475004447.

OpenStudy (anonymous):

correct to finish put it as a percent and your done

OpenStudy (anonymous):

Oh. Thank you :) Have a good night.

OpenStudy (anonymous):

your welcome